A disc rotates about axis its angular velocity w varries with time according to the equation w=at+b initially at t=0 angular velocity 1rad per second and angular position 2rad at t=2sec angular velocity 5rad per sec. Determine angular position and angular acceleration
1
Expert's answer
2019-10-03T09:32:25-0400
"w(0)=b=1\frac{rad}{s}"

The angular acceleration:


"a=\frac{5-1}{2-0}=2\frac{rad}{s^2}"

The angular position:


"\phi=0.5at^2+bt+c=t^2+t+c"

"\phi(0)=c=2\ rad"

Thus,


"\phi=t^2+t+2"